# Perception of Visual Content: Differences Between Humans and Foundation Models

### v1

@@ Abstract
Human-annotated content is often used to train machine learning (ML) models. However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts. This study compares human-generated and ML-generated annotations of images representing diverse socio-economic contexts. We aim to understand differences in perception and identify potential biases in content interpretation. Our dataset comprises images of people from various geographical regions and income levels washing their hands. We compare human and ML-generated annotations semantically and evaluate their impact on predictive models. Our results show low similarity between human and machine annotations from a low-level perspective, i.e., types of words that appear and sentence structures, but are alike in how similar or dissimilar they perceive images across different regions. Additionally, human annotations resulted in best overall and most balanced region classification performance on the class level, while ML Objects and ML Captions performed best for income regression. Humans and machines’ similarity in their lack of bias when perceiving images highlights how they are more alike than what was initially perceived. The superior and fairer performance of using human annotations for region classification and machine annotations for income regression show how important the quality of the images and the discriminative features in the annotations are.
@@ Introduction
Machine Learning (ML) is prevalent in numerous applications, e.g., facial recognition and recommendation systems. However, many of these use cases could potentially result in discrimination and unfair decisions. For example, it is possible that a facial recognition algorithm is less likely to perform well when predicting people of certain ethnic groups, e.g., people of African heritage may not be identified as accurately as people of European origin. This is problematic, especially as we aim to rely more on ML for daily tasks. Since supervised ML models require data for training, biased training data could affect the fairness and bias of the learned model – also known as algorithmic bias – which can cause it to generate prejudiced predictions. Unfortunately, it is quite likely that the data will contain bias considering that they are commonly annotated by human annotators (Demartini, Roitero, and Mizzaro 2024).
Furthermore, generative models have started to be used to produce labelled data for training and evaluation of ML models, which amplifies this issue (Thomas et al. 2023). It is also unclear how these machine annotations are different to human annotations. Since human annotation could contain biases, one possible approach to creating potentially less biased annotations is to use machine-generated annotations. These annotations could likely be more objective compared to those generated by humans.
In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images. This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect fairness in predictive models. This leads to our second research question (RQ2): “How do different combinations of annotations affect fairness in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models. These annotations are based on images of people washing their hands from varying geographical regions and socio-economic classes.
This study contributes to the fields of information retrieval and algorithmic bias in several ways. Firstly, it provides a comparative analysis of human-generated and machine-generated annotations in the context of socio-economic diverse content, which give insights on how humans and machines are alike in their bias when perceiving visual content despite their low-level differences, filling a gap in the existing literature. Secondly, it explores whether different annotations sets can impact the efficacy and fairness of ensemble-based ML models, thus advancing both theoretical and practical understanding of this critical issue.
@@
Effect of Bias on ML Models. It is inevitable that human annotations for training ML models may contain bias and stereotypes. Numerous qualitative and quantitative studies have demonstrated that there exist biases in numerous techniques, datasets, and models related to Natural Language Processing and Information Retrieval (Baeza-Yates 2020; Bolukbasi et al. 2016; Caliskan, Bryson, and Narayanan 2017). These biases could reinforce stereotypes that already exist in society (Rekabsaz, Kopeinik, and Schedl 2021). Fan et al. (2022) investigated this by conducting a study in which participants were asked to annotate video content associated with various regions and levels of income. When sentiment analysis was performed on the annotations, it was found that they contained regional and income-level biases. Unfortunately, as discussed in their paper, one of the main concerns of these biased annotations is that they impact ML models when used as training data. This causes various search and recommendation systems to use these algorithms to feed users with stereotypical content. Additionally, crowd annotators may be experts in different domains, so high-quality annotations are not always guaranteed. Not only could noise be introduced, but the annotated data may not be consistent (Cao et al. 2023). These unreliable annotations may also introduce more biases, which would affect the ML model’s performance. Devillers, Vidrascu, and Lamel (2005) emphasise how these biases can be dangerous when the training dataset involves actual humans as these learned biases tend to mimic forms of discrimination common amongst humans, such as racial and gender bias. An example of this is Microsoft’s “Tay”, an AI chatbot that initially mimicked a teenage girl but became anti-semitic, racist, and sexist over a span of only 16 hours (Zemčík 2020). Sun, Nasraoui, and Shafto (2020) further discuss how the problem worsens when users make decisions based on the biased content provided by these models. These decisions could then result in data used to train other algorithms in the future, thereby creating a never-ending bias-reinforcement cycle. In order to mitigate this issue and break the loop, more diverse data is needed (Fan et al. 2022). Chhikara et al. (2023); Dash et al. (2019) further support this by discussing how textual content written by numerous different authors will contain a greater variation of perspectives, especially on polarising topics, such as religion and politics. This makes it more crucial to see how fairly different perspectives are reflected when attempting to summarise these into one text. van Atteveldt, van der Velden, and Boukes (2021) looked at comparing sentiment analysis methods in news headlines. They used various approaches to compare the sentiment scores produced for these headlines, including manual annotation, crowd-coding, dictionary approaches, and ML algorithms. In this case, their ML algorithm was used to evaluate the news headlines rather than create equivalent annotations to compare against the human annotations. As compared to this body of literature, we focus on comparing human and machine-generated annotations of visual content with the goal of understanding how bias may differ between these annotations.
Dealing with Imbalanced Data. It is important for data to be well-balanced, ensuring that every class in the dataset is well-represented. This is especially true for traditional ML models as they were originally designed for datasets that are balanced. There are various ML methods that were designed to deal with imbalanced datasets, such as random under-sampling and over-sampling techniques (Estabrooks, Jo, and Japkowicz 2004; Chawla et al. 2002), creating synthetic data (Chen et al. 2023; He et al. 2008; Han, Wang, and Mao 2005), and cost-sensitive learning where the model is penalised more when a minority class sample is incorrectly classified compared to a majority class sample (Ting 2002). In our work, we adopt under-sampling techniques to overcome the issue of data imbalance across some classes in our dataset due to the limited number of available samples.
Ensemble Models. Minimising the trade-off between accuracy and fairness is a question that is very relevant to our work. There has been research indicating that a loss in accuracy is inevitable when a model is trained with limitations relating to fairness. Based on Kenfack et al. (2021), we know that compared to individual classifiers, ensemble models have an improved predictive performance, but compromise on fairness. In our work, we include the use of ensemble models in our experiment to test whether there is the same effect on our dataset.
@@
Cognitive Differences in Humans and Machines. To this day, human experts are still needed to provide accurate and reliable annotations that could be used to train ML models (Valizadegan, Nguyen, and Hauskrecht 2013). Understanding how differently humans and machines think is crucial in helping us develop better and fairer automated systems that could assist us in solving problems, i.e., annotating images for ML. Sharifi Noorian et al. (2022) proposed Scalpel-HS, a framework that is intended to identify unknown unknowns in images by combining human intelligence and machine learning. Unknown unknowns refer to instances that an ML model fails to identify or has not been trained on. This framework aims to reduce cognitive load for humans by using ML models to first identify relevant objects and relationships, so that the humans can focus on evaluating relevance rather than performing the task from scratch. In our study, instead of using ML to aid humans in the annotating process, we investigate how similar human annotations and machine-generated annotations are and factors that explain their similarity or dissimilarity considering how humans and machines perceive images.
@@
This section details the methodology we used to address the research questions introduced earlier. The methodology begins with data collection from Dollar Street. Annotations are preprocessed and added to the data either through crowdsourcing or ML techniques. The methodology is divided into two key experiments: the first evaluates annotation similarity; the second investigates the impact of different annotation sets on supervised models, using ensemble learning and comparing the performance of human and machine annotations.
@@
In this study, we utilised CPU resources to conduct all experiments. The computations were performed locally on a device with an Apple M3 Pro Chip, 11 CPU cores, and 36GB of RAM.
@@
The dataset utilised in our study was obtained from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021). Dollar Street visualises global public data by arranging families based on their income levels, from the poorest on the left to the wealthiest on the right. It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes. There are 197 unique instances of images and videos utilised in our experiments, consisting of people washing their hands, sourced from various geographical regions and income levels. The topic of hand washing was selected due to its crucial role in public health and its practice across all cultures and income levels.
@@
Crowdsourcing. The first annotation set is sourced from a structured crowdsourcing process conducted on Amazon Mechanical Turk (MTurk) by Fan et al. (2022). In their study, crowd workers viewed short videos of individuals washing their hands from various global regions and income levels, providing textual annotations. These annotations captured various aspects of handwashing behaviour and the surrounding environmental context. The dataset includes annotations for 112 videos, each annotated by 10 US-based workers. Fan et al. (2022) implemented quality control measures to ensure reliable annotations. Consent from the participants was obtained through a signed consent form. Additionally, no personally identifiable information or offensive content were collected in the crowdsourcing process. We use these existing human annotations in our current study, benefiting from their comprehensive approach to data collection.
ML Object Detection. Images used in this study are retrieved frame-by-frame from videos capturing handwashing practices. The images are processed using a pre-trained Faster R-CNN model with a ResNet-50-FPN backbone, as outlined by Li et al. (2021), which offers improvements over the original model configuration introduced in Ren et al. (2017). This model is available under the MIT License. This model was originally trained on the COCO dataset (Lin et al. 2014) and is utilised to generate our machine-based object annotations, as illustrated in Figure 1. We selected the Faster R-CNN model for its prominence in object detection, its efficiency, and its effectiveness (i.e., mean average precision) on the COCO labels  (Mohammed, Tannouche, and Ounejjar 2022; Ren et al. 2017). We chose to perform object detection on the dataset as most of the images contain objects, i.e., bottles, sinks, instead of subjects performing actions, i.e., person washing hands. Due to this, the COCO dataset was chosen as the most suitable for training as it contains annotations of everyday objects.
@@
The data cleaning process constituted the initial step in preparing our dataset for analysis. In this phase, corrupted images were identified and removed based on criteria including integrity checks and visual inspections. Images that did not contain labels from any of the annotation categories are also removed, i.e., images that did not have any objects detected in it. This process reduced the original dataset to 197 unique data samples with distinct IDs.
The annotations were preprocessed to deal with inconsistencies in the human annotations. These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated.
We utilised a sentence embedding model to convert the three types of annotations into sentence embeddings. Instead of using a pre-trained sentence transformer model to generate the embeddings, we trained a sentence transformer model from scratch on all three sets of annotations. We did this as we want the model and representations to be very focused, and due to the limited number of samples, fine-tuning an existing sentence transformer, such as sBERT, resulted in worse performance.
To create the sentence transformer, we used a pre-trained Transformer model, “distilroberta-base”, which is available at the Hugging Face Hub (Sanh et al. 2019), to obtain the word embedding model. We then combined this word embedding model and a mean pooling function over the token embeddings so that it could generate a fixed-size vector as output. Using the sentence transformer model, we generated embeddings with a size of 384 dimensions.
Experiment 1: Investigating Annotation Similarity (RQ1)
@@
We then used the sentence transformer model’s built-in similarity function to compute the cosine similarity scores between each possible pair of annotations. Using these scores, we look at the average similarity scores for each pair and see how the pairwise scores are distributed using scatterplots and boxplots.
@@
These experiments are appropriate for addressing RQ1 as t-SNE and computation of similarity scores across pairs allows us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
Experiment 2: Exploring How Annotations Impact Fairness in Predictive Models (RQ2)
To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
@@
Before training the models, the dataset was split into training and test sets with a split ratio of 80:20, respectively . The training set was first used to run Grid Search Cross-Validation (CV) with each of the two models to obtain the optimal parameters for each model whilst performing 5-fold cross-validation. Grid Search CV was performed in every independent run.
@@
The first model is an RUS Boost Classifier used to classify each sample on region (Seiffert et al. 2008). This classifier involves the incorporation of the Adaptive Boost algorithm, which is a sequential ensemble technique that can lower a model’s bias and variance, with random under-sampling, which helps deal with data imbalance (Mohammed and Kora 2023; Seiffert et al. 2008). This makes the algorithm ideal for our dataset due to its small size and there not being an equal number of samples from every region. The base estimator used in the model is the Decision Tree Classifier (Swain and Hauska 1977). After obtaining the optimal parameters, i.e., maximum number of estimators where boosting stops and the learning rate, for the model using grid search CV, we trained the classifier using these parameter values. The trained classifier is run 10 times with different random seeds to obtain a mean F1 score for each class (region) and a mean overall weighted F1 score, with their respective standard deviations. This is done to evaluate the model effectiveness for each class and the overall model performance. The model is run multiple times to allow us to evaluate its stability. The Welch’s t-test was performed to compare whether the difference in performance of the best and worst overall-performing models are statistically significant. This test was performed due to the Gaussian distribution of the F1-Scores and varying variances between the models.
@@
The second model uses the Adaboost.R2 algorithm to predict each sample’s income (Drucker 1997). Similar to the region classifier, this algorithm uses a Decision Tree Regressor as its base estimator. Since boosting helps reduce prediction error on the small dataset, this algorithm is suitable for this task. Using grid search CV, we optimised for the number of estimators, learning rate, and the type of loss function used in the Adaboost.R2 algorithm. The model was executed 10 times with different random seeds. The root mean square error (RMSE) was obtained for every run, both for overall and region-based evaluation. The RMSE takes the square root of the mean square error, which performs the measurement using the same units as the target variable, i.e., income level. A regression plot is generated to visualise the distribution of predicted samples by income for the best-performing model. The Kolmogorov-Smirnov test was performed to compare whether the difference in performance of the best and worst overall-performing regression models are statistically significant. This test was performed due to the non-Gaussian distribution of the RMSE scores between the models.
@@
These predictive models are appropriate in addressing RQ2 as it helps us investigate how often each region is correctly identified for a given image, and how well income is predicted for images in a given region. For evaluation purposes, the F1-score is appropriate and sufficient as it is a good metric for evaluating classification performance on imbalanced datasets, as ours is, whereas the RMSE scores are appropriate and sufficient as the measurement being done in the same units as the target variable makes it easier to interpret. Looking at these evaluation metrics for every region helps us determine whether the models perform better for particular regions, giving insight on how fair the model is.
@@
ML Objects and ML Captions 0.87 ± 0.05
ML Objects and Human Labels 0.42 ± 0.1
ML Captions and Human Labels 0.45 ± 0.06
@@
Figure 3: A plot of pairwise relationships between each pair’s similarity score.
Figure 2 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample. The results show how, despite having trained an embedding model from scratch, the embeddings still form into three distinct clusters. Further analysis shows that the human labels cluster is furthest from the ML captions cluster, with the ML object labels cluster in between them.
The clusters of embeddings are due to differences in the annotations. Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images and videos. On the other hand, the ML-generated object labels are primarily object-centric. The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification. The ML-generated captions, although more descriptive than the object labels, tend to not be as specific as the human annotations since they are more likely to contain easily recognisable actions and objects rather than complex descriptions of people’s emotions or the environment. Additionally, there are apparent syntactic differences between the annotations - the ML object labels consists of lists of objects per image; the ML captions consist of a phrase per image; while the human labels are a list of words and phrases. It is also important to note that a limitation that could have affected this distinct categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information. This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations. Considering these differences, the distinct clusters in the visualisation are expected.

### v2

@@ Abstract
Human-annotated content is often used to train machine learning (ML) models. However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts. This study compares human-generated and ML-generated annotations of images representing diverse socio-economic contexts. We aim to understand differences in perception and identify potential biases in content interpretation. Our dataset comprises images of people from various geographical regions and income levels, covering various daily activities and home environments. We compare human and ML-generated annotations semantically and evaluate their impact on predictive models. Our results show highest similarity between ML captions and human labels from a low-level perspective, i.e., types of words that appear and sentence structures, but all three annotations are alike in how similar or dissimilar they perceive images across different regions. Additionally, ML Captions resulted in best overall region classification performance, while ML Objects and ML Captions performed best overall for income regression. The varying performance of annotation sets highlights the notion that all annotations are important, and that human-generated annotations are yet to be replaceable.
@@ Introduction
Data annotations are needed to train Machine Learning (ML) models. These annotations are usually obtained by having human experts manually label data. However, ML models can also be used to produce data annotations, ensuring a more efficient process. However, it is unclear what the differences of such annotations would be as compared to human-generated ones. Using ML models for data annotation could potentially be problematic if the annotations prove to be unreliable or biased. Biased training data could affect the performance and reliability of the learned model – also known as algorithmic bias – which can cause it to generate prejudiced predictions. Unfortunately, it is quite likely that the data will contain bias considering that they are commonly annotated by human annotators (Demartini, Roitero, and Mizzaro 2024). Furthermore, generative models have started to be used to produce labelled data for training and evaluation of ML models, which amplifies this issue (Thomas et al. 2023). It is also unclear how these machine annotations are different to human annotations. Since human annotation could contain biases, one possible approach to creating potentially less biased annotations is to use machine-generated annotations. These annotations could likely be more objective compared to those generated by humans.
In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images. This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect unbiasedness in predictive models. This leads to our second research question (RQ2): “How do different combinations of annotations affect the performance and bias in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models. These annotations are based on images of people performing different activities from varying geographical regions and socio-economic classes.
This study contributes to the fields of information retrieval and algorithmic bias in several ways. Firstly, it presents a new large-scale crowdsourced annotation dataset for data that captures diverse socio-economic visual content. Secondly, it provides a comparative analysis of human-generated and machine-generated annotations in the context of socio-economic diverse content, which gives insights on how humans and machines compare in their bias when perceiving visual content, filling a gap in the existing literature. Thirdly, it explores whether different annotations sets can impact the efficacy and bias of predictive ML models, thus advancing both theoretical and practical understanding of this critical issue.
@@
Effect of Bias on ML Models. It is inevitable that human annotations for training ML models may contain bias and stereotypes. Numerous qualitative and quantitative studies have demonstrated that there exist biases in numerous techniques, datasets, and models related to Natural Language Processing and Information Retrieval (Baeza-Yates 2020; Bolukbasi et al. 2016; Caliskan, Bryson, and Narayanan 2017). These biases could reinforce stereotypes that already exist in society (Rekabsaz, Kopeinik, and Schedl 2021). Fan et al. (2022) investigated this by conducting a study in which participants were asked to annotate video content associated with various regions and levels of income. When sentiment analysis was performed on the annotations, it was found that they contained regional and income-level biases. Unfortunately, as discussed in their paper, one of the main concerns of these biased annotations is that they impact ML models when used as training data. This causes various search and recommendation systems to use these algorithms to feed users with stereotypical content. Additionally, crowd annotators may be experts in different domains, so high-quality annotations are not always guaranteed. Not only could noise be introduced, but the annotated data may not be consistent (Cao et al. 2023). These unreliable annotations may also introduce more biases, which would affect the ML model’s performance. Devillers, Vidrascu, and Lamel (2005) emphasise how these biases can be dangerous when the training dataset involves actual humans as these learned biases tend to mimic forms of discrimination common amongst humans, such as racial and gender bias. Sun, Nasraoui, and Shafto (2020) further discuss how the problem worsens when users make decisions based on the biased content provided by these models. These decisions could then result in data used to train other algorithms in the future, thereby creating a never-ending bias-reinforcement cycle. In order to mitigate this issue and break the loop, more diverse data is needed (Fan et al. 2022). Chhikara et al. (2023); Dash et al. (2019) further support this by discussing how textual content written by numerous different authors will contain a greater variation of perspectives, especially on polarising topics, such as religion and politics. This makes it more crucial to see how fairly different perspectives are reflected when attempting to summarise these into one text. van Atteveldt, van der Velden, and Boukes (2021) looked at comparing sentiment analysis methods in news headlines. They used various approaches to compare the sentiment scores produced for these headlines, including manual annotation, crowd-coding, dictionary approaches, and ML algorithms. In this case, their ML algorithm was used to evaluate the news headlines rather than create equivalent annotations to compare against the human annotations. As compared to this body of literature, we focus on comparing human and machine-generated annotations of visual content with the goal of understanding how bias may differ between these annotations.
Image Captioning. Image captioning has emerged as a critical intersection of computer vision and natural language processing, evolving from template-based approaches to sophisticated deep learning models. As Sharma and Padha (2023) outline, this field has found applications ranging from visual aids to medical imaging and security systems, while highlighting ongoing challenges in real-time performance and domain-specific datasets. Ghandi, Pourreza, and Mahyar (2023) identify key technical challenges including object hallucination, missing context, and illumination conditions, while Xu et al. (2023) frame the captioning process through the stages of seeing, focussing, and telling inspired by humans. Kasai et al. (2021)’s THumB protocol shows that the quality of captions created by humans substantially surpass those generated by machines in terms of covering key information. Moreover, Sarhan and Hegelich (2023) expose critical concerns about AI-generated captions’ tendency toward generic descriptions, which can either erase or stereotype social group identities, particularly in politically sensitive contexts. Building on this literature, our work employs the state-of-the-art BLIP model (Li et al. 2022) with the Vision Transformer architecture (Dosovitskiy et al. 2020), which has shown to be effective among Vision-Language Pre-training methods, to generate captions that we systematically compare with human annotations across socioeconomic contexts.
@@
Cognitive Differences in Humans and Machines. To this day, human experts are still needed to provide accurate and reliable annotations that could be used to train ML models (Valizadegan, Nguyen, and Hauskrecht 2013). Understanding how differently humans and machines think is crucial in helping us develop better and less biased automated systems that could assist us in solving problems, i.e., annotating images for ML. Sharifi Noorian et al. (2022) proposed Scalpel-HS, a framework that is intended to identify unknown unknowns in images by combining human intelligence and machine learning. Unknown unknowns refer to instances that an ML model fails to identify or has not been trained on. This framework aims to reduce cognitive load for humans by using ML models to first identify relevant objects and relationships, so that the humans can focus on evaluating relevance rather than performing the task from scratch. In our study, instead of using ML to aid humans in the annotating process, we investigate how similar human annotations and machine-generated annotations are and factors that explain their similarity or dissimilarity considering how humans and machines perceive images.
@@
In this study, we utilised cloud-based computing resources provided by Amazon Web Services. The experiments used a mix of instance types, with some CPU-only instances, such as the ml.c6i.16xlarge, and some GPU-enabled instances, such as the ml.r5.16xlarge, and ml.r5.24xlarge instances, to accelerate processing times.
@@
The dataset utilised in our study was obtained from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021). Dollar Street visualises global public data by arranging families based on their income levels, from the poorest on the left to the wealthiest on the right. It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes. The dataset contains 3,041 images, sourced from various geographical regions and income levels, consisting of the following image categories: Drinking Water, Drying Clothes, Front Doors, Hand Washing, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes. These images include a diverse range of topics, covering both actions and static elements, such as home environments. This is to provide a good representation of different aspects of daily life, which is applicable across all cultures and income levels.
@@
ML Object Detection. Images used in this study are retrieved frame-by-frame from videos capturing handwashing practices. The images are processed using a pre-trained Faster R-CNN model with a ResNet-50-FPN backbone, as outlined by Li et al. (2021), which offers improvements over the original model configuration introduced in Ren et al. (2017). This model is available under the MIT License. This model was originally trained on the COCO dataset (Lin et al. 2014) and is utilised to generate our machine-based object annotations, as illustrated in Figure 1. We selected the Faster R-CNN model for its prominence in object detection, its efficiency, and its effectiveness (i.e., mean average precision) on the COCO labels  (Mohammed, Tannouche, and Ounejjar 2022; Ren et al. 2017). We chose to perform object detection on the dataset as most of the images contain objects, e.g., bottles, sinks, instead of subjects performing actions, e.g., person washing hands, despite the fact that some of the image categories are meant to be action-based. Due to this, the COCO dataset was chosen as the most suitable for training as it contains annotations of everyday objects.
@@
Figure 2: Screenshot of the MTurk annotation interface showing the task instructions and image assessment layout.
Crowdsourcing. Our human annotations were collected through crowdsourcing tasks on Amazon Mechanical Turk (MTurk). The dataset comprises annotations for 1,886 Dollar Street images covering eight aspects of daily living: drinking water, washing clothes, front doors, hand washing, kitchens, living rooms, places for dinner and washing clothes. We recruited MTurk workers to provide annotations for the images. To ensure high-quality annotations, we restricted participation to workers with at least 5,000 completed tasks and 98% approval rate. Task instructions and interface screenshots are shown in Figure 2. Based on our pilot study that measured average task completion times, participants earned an estimated hourly wage of USD 7.5. Each image was annotated by three different workers to ensure annotation reliability. The total participant compensation for annotating the complete dataset amounted to USD \$373. This research was conducted under the approval of Human Research Ethics Committee. Participants were informed of minimal risks associated with viewing everyday living condition images and provided informed consent before participation. To protect participant privacy and ensure data security, any potentially identifying information was removed from the dataset. We also applied post-processing filters to remove low-quality submissions such as random numbers and nonsensical text. This approach allowed us to build a comprehensive dataset of human observations across Dollar Street images, which represent a wide range of living conditions worldwide.
@@
Prior to obtaining all three annotation sets, the images were selected based on whether they successfully led to the generation of ML Objects with at least 50% confidence and ML Captions beforehand. Moreover, duplicates were removed and down-sampling was performed to ensure the images were balanced for all regions and income groups. The human annotations were processed to deal with inconsistencies. These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated. The number of images in the dataset decreased at each step, starting at 3,041 images at the start, reducing to 2,797 after generating the ML Objects and Captions, then to 1,888 after removing duplicates and down-sampling, and finally to 1,886 unique samples after crowdsourcing. In total, 1,155 samples were discarded from the original dataset.
We utilised a fine-tuned sentence transformer model to convert the three types of annotations into sentence embeddings. We opted to fine-tune an existing model instead of using an pre-trained model out of the box as we want the model and representations to be very focused to our data. We fine-tuned a pre-trained Sentence Transformer model, “all-MiniLM-L12-v2”, available at the Hugging Face Hub [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2. The fine-tuning process was run over 5 epochs with a batch size of 16. The model was fine-tuned on the three annotation sets so that it can better capture the characteristics of the annotations, hence generating closer embeddings. The generated embeddings have a size of 384 dimensions, which is the default embedding size from the original pre-trained model.
Exp. 1: Investigating Annotation Similarity (RQ1)
@@
We then used the sentence transformer model’s built-in cosine similarity function to compute the similarity scores between each possible pair of vector embeddings representing the annotations. Using these scores, we look at the average similarity scores for each pair and see how the pairwise scores are distributed using scatterplots and boxplots.
We computed descriptive statistics about the length of the annotations to help us better understand differences between the sets. We compared the average number of words each worker inputted during crowdsourcing to the average number of words in each caption, as well as the average number of object labels predicted in each image.
@@
These experiments are appropriate for addressing RQ1 as t-SNE, computation of similarity scores across pairs, and annotation length allow us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
Exp. 2: Exploring How Annotations Impact Performance and Bias in Predictive Models (RQ2)
This section emphasises the diversity of model prediction outcomes over overall model effectiveness. These experiments aim to evaluate the distribution of prediction errors across different regions and categories, allowing us to identify potential discrimination, e.g., higher prediction errors for certain subsets of the population as compared to the error on others. To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
@@
As previously mentioned, the dataset was balanced for all geographical regions and income groups, as they are the target variables for these predictive tasks. Since ‘image category’ is not a feature in the training data (annotation embeddings) and is only used in a group-by operation after model training, the metric scores for each category are independent of whether ‘image category’ was balanced. Before training the models, the dataset was split into training and test sets with a split ratio of 80:20, respectively. The training set was first used to run Grid Search Cross-Validation (CV) with each of the two models to obtain the optimal parameters for each model whilst performing 5-fold cross-validation. Grid Search CV was performed in every independent run.
@@
The first model is an RUS Boost Classifier used to classify annotation embeddings on geographical regions, which originate from the dataset, namely Africa, Asia, Europe, and the Americas (Seiffert et al. 2008). This classifier involves the incorporation of the Adaptive Boost algorithm, which is a sequential ensemble technique that can lower a model’s bias and variance (Mohammed and Kora 2023; Seiffert et al. 2008). The base estimator used in the model is the Decision Tree Classifier (Swain and Hauska 1977). After obtaining the optimal parameters, i.e., maximum number of estimators where boosting stops and the learning rate, for the model using grid search CV, we trained the classifier using these parameter values. The trained classifier is run 10 times with different random seeds to obtain a mean F1 score for each class (region) and a mean overall weighted F1 score, with their respective standard deviations. This is done to evaluate the overall model performance, as well as its performance when grouped by geographical region and image category. The model is run multiple times to allow us to evaluate its stability. The Welch’s t-test was performed to compare whether the difference in performance of the best and worst overall-performing models is statistically significant. This test was performed due to the Gaussian distribution of the F1-Scores and varying variances between the models.
@@
The second model uses the Adaboost.R2 algorithm to predict income based on the annotation embeddings (Drucker 1997). Income was chosen as a target variable as it is one of the labels associated with each image in the dataset. Similar to the region classifier, this algorithm uses a Decision Tree Regressor as its base estimator. Since boosting helps reduce prediction error, this algorithm is ideal for this task. Using grid search CV, we optimised for the number of estimators, learning rate, and the type of loss function used in the Adaboost.R2 algorithm. The model was executed 10 times with different random seeds. The root mean square error (RMSE) was obtained for every run to allow for evaluation of the model performance, including overall, by region, and by category. A regression plot was generated to visualise the distribution of predicted samples by income for the best-performing model. The Kolmogorov-Smirnov test was performed to compare whether the difference in performance of the best and worst overall-performing regression models are statistically significant. This test was performed due to the non-Gaussian distribution of the RMSE scores between the models.
@@
These predictive models are appropriate in addressing RQ2 as they help us investigate how well region and income are correctly identified for a given image annotation set. For evaluation, the F1-score is appropriate and sufficient ideal as it is a robust metric which balances precision and recall. The RMSE scores are appropriate and sufficient as the measurement being done in the same units as the target variable makes it easier to interpret. Looking at these evaluation metrics for every region and category helps us determine whether the models perform better for particular regions and categories, giving insight on how biased the model is.
@@
Figure 3 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample. The results show how, despite having a fine-tuned embedding model, the embeddings form into two distinct clusters, with one for ML Captions and Human labels and another for ML Objects. This evidence implies that there exist differences between the different types of annotations we collected in the study. The clusters of embeddings are due to differences in the annotations. Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images. On the other hand, the ML-generated object labels are primarily object-centric. The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification. The ML-generated captions would be more descriptive than the object labels and although they may not be as specific as the human annotations, they are more likely to contain actions and objects, which may also be present in Human Labels. Additionally, there are apparent syntactic differences between the annotations - the ML Objects consists of lists of objects per image; the ML Captions consist of a phrase per image; while the Human Labels are a list of words and phrases. It is also important to note that a limitation that could have affected this categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information. This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations. Considering these differences, the clusters in the visualisation are expected.
@@
ML Objects and ML Captions 0.50 ± 0.12
ML Objects and Human Labels 0.49 ± 0.12
ML Captions and Human Labels 0.69 ± 0.12